The cost per student to attend a school trip is $10.00. At this price, 2000 students will go on the trip. A survey indicates that for every $1 they would lose 100 students. What price per student would result in the greatest revenue.


1
Expert's answer
2021-06-07T16:20:07-0400

Let,

cost per student=p

number of students=n


then,

"TR=p \times n"

At p=10, n=2000; "TR=10\times2000=20,000"

At p=11, n=1900; "TR=11\times 1900=20,900"

At p=12, n=1800; "TR=12\times 1800=21,600"

At p=13, n=1700; "TR=13\times 1700=22,100"

At p=14, n=1600; "TR=14\times 1600=22,400"

At p=15, n=1500; "TR=15\times 1500=22,500"

At p=16, n=1400; "TR=16\times 1400=22,400"

the total revenue will increase with increase in cost per student as the number of students decrease.

Therefore, a price of $15.00 per student would result in the greatest revenue.